OpenStudy (anonymous):

can any help me with this 2(6z-1)-2(z+9)=3(z+1) sorry i typed it wrong before

OpenStudy (mertsj):

You can do this. Start by simplifying the expression in the first parentheses

OpenStudy (mertsj):

What do you get?

OpenStudy (anonymous):

118 - 2z - 18 = 3z + 3 5z = 100 - 3 z = 97/5..

mathslover (mathslover):

\[\huge{120-2-2z-18=3z+3}\] \[\huge{100-2z=3z+3}\] \[\huge{100-3=5z}\] \[\huge{97=5z}\] \[\huge{z=\frac{97}{5}}\]

mathslover (mathslover):

\[\huge{2(6z-1)-2(z+9)=3(z+1)}\] \[\huge{12z-2-2z-18=3z+3}\] \[\huge{10z-20=3z+3}\] \[\huge{10z-3z=23}\] \[\huge{7z=23}\] \[\huge{z=\frac{23}{z}}\]

mathslover (mathslover):

oops sorry i meant \(\huge{z=\frac{23}{7}}\)